# Is There Ethics in Mathematics?

### v2

@@
Intrigued by Borovik’s remarks on the Cambridge University Ethics in Mathematics Project (EiM) in his recent article The Tool/Weapon Duality of Mathematics, revisited [Selected Passages from Correspondence with Friends 13 no. 2 (2025), 15–31. http://bit.ly/4l5LO5z], I decided to find out more about the work of EiM. There is much to read. It seems that a good place to start a discussion with the document titled Four Levels of Ethical Engagement [EiM Discussion Paper 1/2018 University of Cambridge Ethics in Mathematics Project, https://www.ethics-in-mathematics.com/assets/dp/18_1.pdf] by Maurice Chiodo and Piers Bursill-Hall. Unless indicated otherwise, all quotes in my text are from this paper. In the introduction, the authors state
@@ 1. Who are those mathematicians?
Do you know such mathematicians? I am sure you do, and so do I. Is this a “reasonable generalisation of the norms of mathematicians’ behaviour.”I doubt. And one more quote from the final section of the paper:
@@
At that time, all exams had both written and oral components. Oral exams were given by the professors teaching the course, and some of those professors were legends of the Polish mathematics. I can say with confidence that we got to know all of them relatively well. I do not remember a single person that would fit the stereotype described by Chiodo and Bursill. They were very divers as individuals, but at the same time quite united by their attitude towards mathematics as a subject and by the understanding of its role in the society.
@@ 2. Engagement
In 1980, the Solidarity movement became an official resistance force until it was crushed by the imposition of the martial law in December of 1981. This eventually led to the collapse of the system, but only in 1989. Many mathematicians were active agents of those developments. I will mention two of thenm.
@@
The list of mathematical contributions includes papers by Max Dickman, Per Lindström, Peter Aczel, Jeff Paris, and Maurice Boffa—mathematicians I got to know personally. In my early years in mathematics, they were my role models. I learned mathematics from them, but I also paid much attention to their high ethical standards. There were many more who either visited Poland at that time or I had a chance to meet at conferences and who impressed me very much. As I think now about those days, more and more names come to mind of mathematicians from all over Europe and the US I got to know. All of them open mined and actively involved in politics at various levels. I will mention just two more names. The full list would be very long.
@@ 3. Ethics and education
I absolutely agree that communication is a problem. I have been there, I have done that. During my career I had a chance to speak publicly at meetings of various levels of importance at my university. I am not a bad speaker and I have a reputation of being a good teacher. I am willing and able to explain mathematics to non-mathematicians; however, there are limits to what one can do, and that is because the non-mathematicians one talks to often know very little about mathematics, and this is not necessarily their fault.333 A curious fact: In my academic discussions, I learned that it is actually sometimes easier to talk to our colleagues in the humanities than in the disciplines that explicitly use mathematics.

### v3

@@
Intrigued by Borovik’s remarks on the Cambridge University Ethics in Mathematics Project Mathematics Project (EiM) in his recent article The Tool/Weapon Duality of Mathematics, revisited [Selected Passages from Correspondence with Friends 13 no. 2 (2025), 15–31. http://bit.ly/4l5LO5z], I decided to find out more about the work of EiM. There is much to read. It seems that a good place to start a discussion with the document titled Four Levels of Ethical Engagement [EiM Discussion Paper 1/2018 University of Cambridge Ethics in Mathematics Project, https://www.ethics-in-mathematics.com/assets/dp/18_1.pdf] by Maurice Chiodo and Piers Bursill-Hall. Unless indicated otherwise, all quotes in my text are from this paper. In the introduction, the authors state
@@ 1. Who are those mathematicians?
Do you know such mathematicians? I am sure you do, and so do I. Is this a “reasonable generalisation of the norms of mathematicians’ behaviour”? I doubt. And one more quote from the final section of the paper:
@@
At that time, all exams had both written and oral components. Oral exams were given by the professors teaching the course, and some of those professors were legends of the Polish mathematics. I can say with confidence that we got to know all of them relatively well. I do not remember a single person that would fit the stereotype described by Chiodo and Bursill. They were a diverse group, but at the same time quite united by their attitude towards mathematics as a subject and by the understanding of its role in the society.
@@ 2. Engagement
In 1980, the Solidarity movement became an official resistance force until it was crushed by the imposition of the martial law in December of 1981. This eventually led to the collapse of the system, but only in 1989. Many mathematicians were active agents of those developments. I will mention two of them.
@@
The list of mathematical contributions includes papers by Max Dickmann, Per Lindström, Peter Aczel, Jeff Paris, and Maurice Boffa—mathematicians I got to know personally. In my early years in mathematics, they were my role models. I learned mathematics from them, but I also paid much attention to their high ethical standards. There were many more who either visited Poland at that time or I had a chance to meet at conferences and who impressed me very much. As I think now about those days, more and more names come to mind of mathematicians from all over Europe and the US I got to know. All of them open mined and actively involved in politics at various levels. I will mention just two more names. The full list would be very long.
@@ 3. Ethics and education
I absolutely agree that communication is a problem. I have been there, I have done that. During my career I had a chance to speak publicly at meetings of various levels of importance at my university. I am not a bad speaker and I have a reputation of being a good teacher. I am willing and able to explain mathematics to non-mathematicians; however, there are limits to what one can do, and that is because the non-mathematicians one talks to often know very little about mathematics, and this is not necessarily their fault. [^3] A curious fact: In my academic discussions, I learned that it is actually sometimes easier to talk to our colleagues in the humanities than in the disciplines that explicitly use mathematics.